# Prospects for detection of ultra high frequency gravitational waves from compact binary coalescenses with resonant cavities

### v2

@@ III.1.3 Effective time limited by the sampling rate
Even if this case seems to be phenomenologicaly more interesting because it corresponds to chirp masses ${\cal M}_{c}>\mathcal{M}_{\text{max}}^{\Delta\nu}\sim 10^{-12}M_{\odot}$, we remind the reader that the chirp mass can, anyway, not be too high, otherwise the black holes would be located inside their ISCO (or would even already have merged), as shown on Fig. 1.
@@
The values for the other setups are: $8.1\times 10^{9}$ for the $0.34$ GHz configuration and $1.8\times 10^{8}$ for the $11.47$ GHz configuration.
III.2 Lock-in amplifier – monochromatic sources
Let us now consider the case corresponding to a sufficiently monochromatic source, detected by a phase-sensitive instrument such as a lock-in amplifier. In the SNR formula, the bandwidth then has to be taken at the frequency resolution, that is $1/t_{\text{int}}$ Berlin et al. (2022a) where $t_{\text{int}}$ is the integration time. The physical criterion for this to be possible is that the frequency drift during the time $t_{\text{int}}$ is much smaller than $\Delta\nu$, which reads
$\dot{f}t_{\text{int}}\ll\frac{\nu}{Q}\Leftrightarrow t_{\text{int}}\ll t_{\Delta\nu}.$ (21)
As expected, this corresponds to the case where the measurement is limited by the duration of the experiment.
The signal to noise ratio now reads Berlin et al. (2022a):
@@ III.2 Lock-in amplifier – monochromatic sources
$h>\sqrt{2\mu_{0}c^{2}k_{B}}(2\pi\nu)^{-\frac{3}{2}}\eta^{-1}B_{0}^{-1}V_{\text{cav}}^{-\frac{5}{6}}Q^{-\frac{1}{2}}T_{\text{sys}}^{\frac{1}{2}}t_{\text{int}}^{-\frac{1}{2}},$ (23)
Which translates in
$D$ $<$ $4\pi^{\frac{13}{6}}c^{-5}G^{\frac{5}{3}}\mu_{0}^{-\frac{1}{2}}k_{B}^{-\frac{1}{2}}\nu^{\frac{13}{6}}\eta B_{0}V_{\text{cav}}^{\frac{5}{6}}Q^{\frac{1}{2}}T_{\text{sys}}^{-\frac{1}{2}}{\cal M}_{c}^{\frac{5}{3}}t_{\text{int}}^{\frac{1}{2}}.$ (24)
It should be noticed that the exponents are quite different from those appearing in the previous case, making the functional dependence quite interesting.
Plugging $t_{\text{int}}=(1/100)t_{\Delta\nu}$ leads to
$h$ $>$ $4\sqrt{30}\pi^{-\frac{1}{6}}\mu_{0}^{\frac{1}{2}}k_{B}^{\frac{1}{2}}c^{-\frac{3}{2}}G^{\frac{5}{6}}\nu^{-\frac{1}{6}}\eta^{-1}B_{0}^{-1}V_{\text{cav}}^{-\frac{5}{6}}T_{\text{sys}}^{\frac{1}{2}}{\cal M}_{c}^{\frac{5}{6}}$ (25)
@@
$\Leftrightarrow h$ $>$ $7.0\times 10^{-24}\times\bigg{(}\frac{2.67\ \text{GHz}}{\nu}\bigg{)}^{\frac{1}{6}}\bigg{(}\frac{0.1}{\eta}\bigg{)}\bigg{(}\frac{27\ \text{T}}{B_{0}}\bigg{)}\bigg{(}\frac{1.83\times 10^{-3}\ \text{m}^{3}}{V_{\text{cav}}}\bigg{)}^{\frac{5}{6}}\bigg{(}\frac{T_{\text{sys}}}{0.4\ \text{K}}\bigg{)}^{\frac{1}{2}}\left(\frac{{\cal M}_{c}}{10^{-17}M_{\odot}}\right)^{\frac{5}{6}}~{},$
The prefactors are $2.4\times 10^{-25}$ for the 0.34 GHZ configuration and $1.1\times 10^{-22}$ for the 11.47 GHZ configuration. The distance is
$D$ $<$ $\frac{1}{2\sqrt{30}}\mu_{0}^{-\frac{1}{2}}k_{B}^{-\frac{1}{2}}\pi^{\frac{5}{6}}G^{\frac{5}{6}}c^{-\frac{5}{2}}\nu^{\frac{5}{6}}\eta B_{0}V_{\text{cav}}^{\frac{5}{6}}T_{\text{sys}}^{-\frac{1}{2}}{\cal M}_{c}^{\frac{5}{6}}$ (26)
@@
$\Leftrightarrow D$ $<$ $23\times\left(\frac{\nu}{2.67~{}\text{GHz}}\right)^{\frac{5}{6}}\left(\frac{\eta}{0.1}\right)\left(\frac{B_{0}}{27\ \text{T}}\right)\left(\frac{V_{\text{cav}}}{1.83\times 10^{-3}\ \text{m}^{3}}\right)^{\frac{5}{6}}\left(\frac{0.4~{}\text{K}}{T_{\text{sys}}}\right)^{\frac{1}{2}}\left(\frac{{\cal M}_{c}}{10^{-17}M_{\odot}}\right)^{\frac{5}{6}}~{}\text{m}~{}.$
The prefactors are $1.7\times 10^{2}$ for the 0.34 GHZ configuration and $3.9$ for the 11.47 GHZ configuration. The quality factor does not anymore enters directly the results.
In practice, a numerical filter might be used in this case to go beyond the $\Delta\nu$ of the cavity. We leave the detailed analysis for a future work.
III.3 Summary of the results
Figures 4 and 5 summarize the full study by presenting the strain sensitivity and the maximum distance at which the source can be situated to be detectable as functions of the chirp mass in all the considered regimes. Interestingly, large distances are reached for chirp masses where the instrument is actually less sensitive. This is simply because the signal is way more important in this range and this over-compensates the previous effect.
FIG. 4: Strain sensitivity of the GrAHal 2.67 GHz configuration as a function of the chirp mass of the physical system for $t_{\text{max}}=1$ min. The black dashed line corresponds to $t_{\text{max}}=1$ year. The purple dashed line corresponds to the lock-in case. The hatched region on the right is excluded by the results presented Fig.1.
FIG. 5: Maximum distance to the inspiralling system for the GrAHal 2.67 GHz configuration as a function of the chirp mass of the physical system for $t_{\text{max}}=1$ min. The black dashed line corresponds to $t_{\text{max}}=1$ year. The purple dashed line corresponds to the lock-in case. The hatched region on the right is excluded by the results presented Fig.1.
IV Event rate
We now assume that both black holes have the same mass. This is the optimistic scenario. Binaries may form if the black holes were produced very close one to the other so that their dynamics decoupled from the background expansion early enough. This also requires a specific external gravitational influence to prevent the system from directly collapsing. If all the required conditions are met, and assuming that the density of black holes saturates the upper limit derived from microlensing Niikura et al. (2019a), the merging rate can be shown to be Aggarwal et al. (2021); Raidal et al. (2019); Kocsis et al. (2018); Gow et al. (2020):
$\tau\propto{\cal M}_{c}^{-\frac{32}{37}},$ (27)
or, numerically,
$\tau$$\approx$ $300\bigg{(}\frac{m}{M_{\odot}}\bigg{)}^{-\frac{32}{37}}\text{yr}^{-1}\text{Gpc}^{-3}$ (28)
$\approx$ $1.1\times 10^{-74}\bigg{(}\frac{m}{M_{\odot}}\bigg{)}^{-\frac{32}{37}}\text{yr}^{-1}\text{m}^{-3}$ (29)
$\approx$ $1.3\times 10^{11}\bigg{(}\frac{m}{10^{-10}M_{\odot}}\bigg{)}^{-\frac{32}{37}}\text{yr}^{-1}\text{Gpc}^{-3}$ (30)
$\approx$ $4.8\times 10^{-66}\bigg{(}\frac{m}{10^{-10}M_{\odot}}\bigg{)}^{-\frac{32}{37}}\text{yr}^{-1}\text{m}^{-3},$ (31)
where $m=m_{1}=m_{2}=2^{1/5}{\cal M}_{c}$ (in practice $m\sim{\cal M}_{c}$). The number $N$ of expected events reads $N=4\pi\tau t_{\text{tot}}D^{3}/3$ where $t_{\text{tot}}$ is the total time of the experimental data taking. This later should not to be confused with the previous time scales entering the different estimates and which were basically corresponding to the sampling time, to the drift duration of the signal in the bandwidth of the experiment, or to the duration of a single “run”. Obviously, Eq. (27) is, in principle, suitable on cosmological scales and a substantial boost factor can be expected when dealing with nearby phenomena. This would however not alter the conclusions.
In the case of an integration time limited by the signal drift, the number of expected events reads:
$N\approx 1.8\times 10^{-50}$$\times$ $\left(\frac{\nu}{2.67\ \text{GHz}}\right)^{\frac{15}{4}}\left(\frac{\eta}{0.1}\right)^{3}\left(\frac{B_{0}}{27\ \text{T}}\right)^{3}\left(\frac{V_{\text{cav}}}{1.83\times 10^{-3}\ \text{m}^{3}}\right)^{\frac{5}{2}}\left(\frac{Q}{10^{5}}\right)^{\frac{3}{2}}$

### v3

@@ III.1.3 Effective time limited by $t_{\Delta\nu}$
The values for the other setups are: $8.1\times 10^{9}~{}\text{m}$ for the $0.34$ GHz configuration and $1.8\times 10^{8}~{}\text{m}$ for the $11.47$ GHz configuration.
iii) Cavity charging effect
We now forget the previous effect and we focus on another effect limiting the sensitivity for short signals: the fact that the $Q$ factor entering the expression of the power, Eq. (6), should not be the cavity quality factor but the signal quality factor when the latter is smaller. For masses such that $t_{\Delta\nu}<\Delta\nu^{-1}$, one therefore has to replace: $Q\rightarrow\nu t_{\Delta\nu}=Qt_{\Delta\nu}/t_{\text{min}}$ Kim et al. (2020). This can be taken into account by setting $t_{eff}=t_{\Delta\nu}^{3}/t_{\text{min}}^{2}$, which leads to:
$h$ $>$ $\frac{1}{2}\left(\frac{96}{5}\right)^{\frac{3}{4}}\pi^{\frac{1}{2}}\mu_{0}^{\frac{1}{2}}k_{B}^{\frac{1}{2}}c^{-\frac{11}{4}}G^{\frac{5}{4}}\nu^{\frac{1}{4}}\eta^{-1}B_{0}^{-1}V_{\text{cav}}^{-\frac{5}{6}}Q^{\frac{1}{2}}T_{\text{sys}}^{\frac{1}{2}}{\cal M}_{c}^{\frac{5}{4}}$ (20)
$\Leftrightarrow h$ $>$ $4.3\times 10^{-17}\times\left(\frac{\nu}{2.67~{}\text{GHz}}\right)^{\frac{1}{4}}\left(\frac{0.1}{\eta}\right)\left(\frac{27~{}\text{T}}{B_{0}}\right)\left(\frac{1.83\times 10^{-3}\text{m}^{3}}{V_{\text{cav}}}\right)^{\frac{5}{6}}\left(\frac{Q}{10^{5}}\right)^{\frac{1}{2}}\left(\frac{T_{\text{sys}}}{0.4~{}\text{K}}\right)^{\frac{1}{2}}\left(\frac{{\cal M}_{c}}{10^{-9}M_{\odot}}\right)^{\frac{5}{4}}.$
The numerical prefactors for the other setups are: $6.2\times 10^{-19}$ for the $0.34$ GHz configuration and $1.2\times 10^{-15}$ for the $11.47$ GHz configuration.
The associated maximal distance at which a merger could be observed is thus:
$D$ $<$ $4\bigg{(}\frac{5}{96}\bigg{)}^{\frac{3}{4}}\pi^{\frac{1}{6}}c^{-\frac{5}{4}}G^{\frac{5}{12}}\mu_{0}^{-\frac{1}{2}}k_{B}^{-\frac{1}{2}}\nu^{\frac{5}{12}}\eta B_{0}V_{\text{cav}}^{\frac{5}{6}}Q^{-\frac{1}{2}}T_{\text{sys}}^{-\frac{1}{2}}{\cal M}_{c}^{\frac{5}{12}}$ (21)
$\Leftrightarrow D$ $<$ $8.3\times 10^{7}\times\left(\frac{\nu}{2.67~{}\text{GHz}}\right)^{\frac{5}{12}}\left(\frac{\eta}{0.1}\right)\left(\frac{B_{0}}{27\ \text{T}}\right)\left(\frac{V_{\text{cav}}}{1.83\times 10^{-3}\ \text{m}^{3}}\right)^{\frac{5}{6}}\left(\frac{10^{5}}{Q}\right)^{\frac{1}{2}}\left(\frac{0.4~{}\text{K}}{T_{\text{sys}}}\right)^{\frac{1}{2}}\left(\frac{{\cal M}_{c}}{10^{-9}M_{\odot}}\right)^{\frac{5}{12}}~{}\text{m}.$
The numerical prefactors for the other setups are: $1.5\times 10^{9}~{}\text{m}$ for the $0.34$ GHz configuration and $7.5\times 10^{6}~{}\text{m}$ for the $11.47$ GHz configuration.
iv) Full effect
To get a conservative sensitivity estimate for the ”high masses range” we now take into account both previous effects simultaneously. This can be accounted for by setting $t_{\rm eff}=t_{\Delta\nu}^{4}/t_{\text{min}}^{3}$ (see the appendix for more details and comments on this treatment). Still requiring $SNR>1$ for detection, this leads to a minimum strain for detection of:
$h$ $>$ $\frac{48}{5}\pi^{\frac{7}{6}}c^{-4}G^{\frac{5}{3}}\mu_{0}^{\frac{1}{2}}k_{B}^{\frac{1}{2}}\nu^{\frac{2}{3}}\eta^{-1}B_{0}^{-1}V_{\text{cav}}^{-\frac{5}{6}}QT_{\text{sys}}^{\frac{1}{2}}{\cal M}_{c}^{\frac{5}{3}}$ (22)
$\Leftrightarrow h$ $>$ $5.6\times 10^{-16}\times\bigg{(}\frac{\nu}{2.67\ \text{GHz}}\bigg{)}^{\frac{2}{3}}\bigg{(}\frac{0.1}{\eta}\bigg{)}\bigg{(}\frac{27\ \text{T}}{B_{0}}\bigg{)}\bigg{(}\frac{1.83\times 10^{-3}\ \text{m}^{3}}{V_{\text{cav}}}\bigg{)}^{\frac{5}{6}}\bigg{(}\frac{Q}{10^{5}}\bigg{)}\bigg{(}\frac{T_{\text{sys}}}{0.4\ \text{K}}\bigg{)}^{\frac{1}{2}}\left(\frac{{\cal M}_{c}}{10^{-9}M_{\odot}}\right)^{\frac{5}{3}}~{}.$
@@
The values of the numerical prefactors for the other setups are respectively: $3.4\times 10^{-18}$ for the $0.34$ GHz configuration and $3.0\times 10^{-14}$ for the $11.47$ GHz configuration.
The associated reachable distance becomes:
$D$ $<$ $\frac{5}{24}\pi^{-\frac{1}{2}}\mu_{0}^{-\frac{1}{2}}k_{B}^{-\frac{1}{2}}\eta B_{0}V_{\text{cav}}^{\frac{5}{6}}Q^{-1}T_{\text{sys}}^{-\frac{1}{2}}$ (23)
$\Leftrightarrow D$ $<$ $6.3\times 10^{6}\times\left(\frac{\eta}{0.1}\right)\left(\frac{B_{0}}{27\ \text{T}}\right)\left(\frac{V_{\text{cav}}}{1.83\times 10^{-3}\ \text{m}^{3}}\right)^{\frac{5}{6}}\left(\frac{10^{5}}{Q}\right)\left(\frac{0.4~{}\text{K}}{T_{\text{sys}}}\right)^{\frac{1}{2}}~{}\text{m}.$
The values for the other setups are: $2.6\times 10^{8}~{}\text{m}$ for the $0.34$ GHz configuration and $3.1\times 10^{5}~{}\text{m}$ for the $11.47$ GHz configuration.
III.2 Optimized resolution bandwidths
In cases where the signal is more coherent than the cavity ($\frac{1}{\Delta\nu}<t_{\Delta\nu}\Leftrightarrow\dot{f}t_{\text{int}}<\frac{\nu}{Q}$), phase sensitive detectors allow to optimize the signal-to-noise ratio by setting the resolution bandwidth at the signal spectral width.
i) Effective time limited by the duration of the experiment
In the ”low-mass” regime corresponding to $t_{\Delta\nu}>t_{\text{max}}$ the signal spectral width is given by $\Delta\nu_{\text{signal}}=t_{\text{max}}^{-1}$ and the signal to noise ratio now reads:
@@ III.2 Optimized resolution bandwidths
$h$ $>$ $\frac{1}{2}\pi^{-\frac{3}{2}}c\mu_{0}^{\frac{1}{2}}k_{B}^{\frac{1}{2}}\nu^{-\frac{3}{2}}\eta^{-1}B_{0}^{-1}V_{\text{cav}}^{-\frac{5}{6}}Q^{-\frac{1}{2}}T_{\text{sys}}^{\frac{1}{2}}t_{\text{max}}^{-\frac{1}{2}}$ (25)
$\Leftrightarrow h$ $>$ $1.5\times 10^{-23}\times\bigg{(}\frac{2.67\ \text{GHz}}{\nu}\bigg{)}^{\frac{3}{2}}\bigg{(}\frac{0.1}{\eta}\bigg{)}\bigg{(}\frac{27\ \text{T}}{B_{0}}\bigg{)}\bigg{(}\frac{1.83\times 10^{-3}\ \text{m}^{3}}{V_{\text{cav}}}\bigg{)}^{\frac{5}{6}}\bigg{(}\frac{10^{5}}{Q}\bigg{)}^{\frac{1}{2}}\bigg{(}\frac{T_{\text{sys}}}{0.4\ \text{K}}\bigg{)}^{\frac{1}{2}}\bigg{(}\frac{60\ \text{s}}{t_{\text{max}}}\bigg{)}^{\frac{1}{2}}~{},$
which translates into
$D$ $<$ $4\pi^{\frac{13}{6}}c^{-5}G^{\frac{5}{3}}\mu_{0}^{-\frac{1}{2}}k_{B}^{-\frac{1}{2}}\nu^{\frac{13}{6}}\eta B_{0}V_{\text{cav}}^{\frac{5}{6}}Q^{\frac{1}{2}}T_{\text{sys}}^{-\frac{1}{2}}{\cal M}_{c}^{\frac{5}{3}}t_{\text{max}}^{\frac{1}{2}}$ (26)
@@
$\Leftrightarrow D$ $<$ $11\times\bigg{(}\frac{\nu}{2.67\ \text{GHz}}\bigg{)}^{\frac{13}{6}}\left(\frac{\eta}{0.1}\right)\left(\frac{B_{0}}{27\ \text{T}}\right)\left(\frac{V_{\text{cav}}}{1.83\times 10^{-3}\ \text{m}^{3}}\right)^{\frac{5}{6}}\left(\frac{Q}{10^{5}}\right)^{\frac{1}{2}}\left(\frac{0.4~{}\text{K}}{T_{\text{sys}}}\right)^{\frac{1}{2}}\left(\frac{{\cal M}_{c}}{10^{-17}M_{\odot}}\right)^{\frac{5}{3}}\bigg{(}\frac{t_{\text{max}}}{60\ \text{s}}\bigg{)}^{\frac{1}{2}}~{}\text{m}.$
ii) Effective time limited by the duration of the signal
In the ”intermediate-mass” regime for which $t_{\text{min}}<t_{\Delta\nu}<t_{\text{max}}$ the signal spectral width is given by $\Delta\nu_{\text{signal}}=t_{\Delta\nu}^{-1}$ and the signal to noise ratio now reads:
${\rm SNR}\sim\frac{P_{sig}}{k_{B}T_{\text{sys}}}t_{\Delta\nu}.$ (27)
This leads to
$h$ $>$ $\sqrt{\frac{24}{5}}\pi^{-\frac{1}{6}}c^{-\frac{3}{2}}G^{\frac{5}{6}}\mu_{0}^{\frac{1}{2}}k_{B}^{\frac{1}{2}}\nu^{-\frac{1}{6}}\eta^{-1}B_{0}^{-1}V_{\text{cav}}^{-\frac{5}{6}}T_{\text{sys}}^{\frac{1}{2}}{\cal M}_{c}^{\frac{5}{6}}$ (28)
@@
$\Leftrightarrow h$ $>$ $2.2\times 10^{-22}\times\bigg{(}\frac{2.67\ \text{GHz}}{\nu}\bigg{)}^{\frac{1}{6}}\bigg{(}\frac{0.1}{\eta}\bigg{)}\bigg{(}\frac{27\ \text{T}}{B_{0}}\bigg{)}\bigg{(}\frac{1.83\times 10^{-3}\ \text{m}^{3}}{V_{\text{cav}}}\bigg{)}^{\frac{5}{6}}\bigg{(}\frac{T_{\text{sys}}}{0.4\ \text{K}}\bigg{)}^{\frac{1}{2}}\left(\frac{{\cal M}_{c}}{10^{-14}M_{\odot}}\right)^{\frac{5}{6}}~{},$
which translates into
$D$ $<$ $\sqrt{\frac{5}{6}}\pi^{\frac{5}{6}}c^{\frac{-5}{2}}G^{\frac{5}{6}}\mu_{0}^{-\frac{1}{2}}k_{B}^{-\frac{1}{2}}\nu^{\frac{5}{6}}\eta B_{0}V_{\text{cav}}^{\frac{5}{6}}T_{\text{sys}}^{-\frac{1}{2}}{\cal M}_{c}^{\frac{5}{6}}$ (29)